# Full Line Code Completion: Bringing AI to Desktop

### v1

@@ 4.1. Offline Evaluation
With such a setup, we run primary experiments for every major change implemented in FLCC. For example, while implementing our strategy for collecting hypotheses described in Section 3.4.2, we observed that its usage increased the percentage of Perfect Lines from 13% to 21%, while Matched Ratio only dropped from 20.6% to 20.3%, which is quite negligible.
4.2. Online Evaluation
Offline evaluation is a valuable tool, but it might be biased, so it is not enough to make reliable decisions about product changes. As described in Section 2.1, at JetBrains, we use early access programs (EAPs) to collect anonymous logs for experimental features and evaluate their performance. For ML features, in particular, we pay attention to rigorous A/B testing during EAPs. Such an approach allows to compare user experience with only standard completion versus both standard and Full Line Code Completion enabled. Also, this provides a convenient way of comparing different implementations of FLCC itself. During such an A/B testing for Python in the fall of 2023, we collected a lot of different important metrics from literature (Ziegler et al., 2022; Wang et al., 2023), and here are the general results for them.
Ratio of completed code. This is our main, golden star metric used for the assessment of code completion quality. It is defined as a ratio of symbols of code written with code completion among all the written code. The numerator of the metric includes code from all code completion sources: standard and inline completions, so it allows us to assess the whole code completion experience.
The metric bumped 1.5 times (from 16% to 24%) for users with FLCC enabled compared to users with standard code completion only. The result is statistically significant with a p-value ¡ 0.001.
Acceptance rate. This is the ratio of events where Full Line Completion was selected to events where it was shown. Previous studies (Ziegler et al., 2022) show that the acceptance rate strongly correlates with the perceived quality of code completion.
The acceptance rate for inline completion was about 15%. For standard completion, the value is about 27%, however, no conclusions can be made from this comparison due to significantly different lengths of suggestions. Our solution does not perform as well as GitHub Copilot, which shows an acceptance rate of about 29% but uses a model that is two magnitudes larger and is cloud-based.
Persistence rate. This metric can be defined differently (Ziegler et al., 2022), and it accounts for the ratio of code that stays in the file for some time after accepting code completion. In our approach, we track the average number of cancellation actions, including Ctrl/Cmd+Z shortcuts, and various text deletion events after the code completion suggestion was selected.
@@ 4.2. Online Evaluation
Time to show. This metric calculates CPU time from Full Line Completion invocation to the showing of a suggestion.
@@ 5. Discussion
Challenges and future directions. Despite the evident advantages, implementing FLCC in a local environment presented unique challenges, especially considering the constraints of memory footprint and processing speed. While we overcame these challenges, they highlight the need for optimizing machine learning models for practical use in constrained environments.
Additionally, since the feature is currently focused on Python in PyCharm Pro and DataSpell, there is significant scope for expanding its capabilities to other programming languages and IDEs. This expansion could not only increase the user base but also offer insights into how different programming languages and development environments interact with AI-driven code completion tools.
@@ 6. Threats to Validity
Limitations of offline evaluation. We use a limited number of repositories and files to run our offline evaluation pipeline. Despite our effort to collect repositories with diverse code, our dataset for offline evaluation remains limited in terms of size and might not represent all Python code well enough.
Online evaluation bias. Our study utilized anonymous data from Early Access Program (EAP) users to evaluate various code completion approaches via A/B testing. Although the results were statistically significant, they may not fully represent the broader user base of PyCharm Pro or DataSpell, as EAP participants typically have a deeper understanding and engagement with the IDE’s features, and their usage patterns may differ from average users.
@@
7. Conclusion
In this work, we shared our approach to implementing a multi-token code completion feature for IntelliJ-based IDEs, called Full Line Code Completion. The feature fills a niche of local-based multi-token code completion engines, being available for a wide range of users, including those with limited internet access and those with privacy concerns. It meets the speed and memory footprint restrictions, as well as fulfills several other important design principles. The resulting implementation has proven its efficiency during A/B testing on real users and was bundled into PyCharm Pro and DataSpell 2023.3 release versions. Our ongoing and further work is focused on extending the feature to a broader range of languages and IDEs, as well as improving the quality for existing ones. The latter is going to be primarily executed via state-of-the-art language models like LLaMA. We hope that our work can prove useful to both researchers and practitioners, and showcase the lifecycle of a complex research project in the real product.

### v2

@@ IV-A Offline Evaluation
With such a setup, we run primary experiments for every major change implemented in FLCC. For example, while implementing our strategy for collecting hypotheses described in Section III-D2, for Python, we observed that its usage increased the percentage of Perfect Lines from 13% to 21%, while Matched Ratio only dropped from 20.6% to 20.3%, which is quite negligible.
IV-B Online Evaluation
Offline evaluation is a valuable and convenient tool, but it might be biased, so it is not enough to make reliable decisions about product changes. As described in Section II-A, at JetBrains, we use early access programs (EAPs) to collect anonymous logs for experimental features and evaluate their performance. For ML features, in particular, we pay attention to rigorous A/B testing during EAPs. Such an approach allows to compare user experience with only standard completion versus both standard and Full Line Code Completion enabled. Also, this provides a convenient way of comparing different implementations of FLCC itself. During such A/B testing in late 2023 and 2024, we collected a lot of data to compute different important metrics from literature [62, 63]. We share the overall comparison for multiple languages in Table I and describe the metrics below, highlighting results for Python.
TABLE I: The results of online evaluation for different languages.
Language Ratio of completed code Acceptance rate
without FLCC with FLCC with FLCC
[…]
Ratio of completed code. This is our main, golden star metric used for the assessment of code completion quality. It is defined as a ratio of symbols of code written with code completion among all the written code. The numerator of the metric includes code from all code completion sources: standard and inline completions, so it allows us to assess the entire code completion experience.
For Python, the metric bumped 1.3 times (from 17% to 22%) for users with FLCC enabled compared to users with standard code completion only. The result is statistically significant with a p-value $<$ 0.001.
Acceptance rate. This is the ratio of events where Full Line Code Completion was selected to events where it was shown. Previous studies [62] show that the acceptance rate strongly correlates with the perceived quality of code completion.
The acceptance rate for Full Line Code Completion for Python was about 38%. For standard completion, the value is about 27%, however, no conclusions can be made from this comparison due to significantly different lengths of suggestions. Our solution does not perform as well as GitHub Copilot, which shows an acceptance rate of about 40% for both single-line and multi-line suggestions combined. However, one should note that GitHub Copilot uses a model that is two magnitudes larger and is cloud-based.
Persistence rate. This metric can be defined differently [62], and it accounts for the ratio of code that stays in the file for some time after accepting code completion. In our approach, we track the average number of cancellation actions, including Ctrl/Cmd+Z shortcuts, and various text deletion events after the code completion suggestion was selected.
@@ IV-B Online Evaluation
Time to show. This metric calculates CPU time from the invocation of completion to the showing of a suggestion.
@@ V Discussion
Challenges and future directions. Despite the evident advantages, implementing FLCC in a local environment presented unique challenges, especially considering the constraints of memory footprint and processing speed. While we overcame these challenges, they highlight the need for optimizing ML models for practical use in constrained environments.
Additionally, the feature’s performance is significantly different for different languages, so we are aiming to study these differences and gain more insights about code completion for all those languages.
@@ VI Threats to Validity
Limitations of offline evaluation. We use a limited number of repositories and files to run our offline evaluation pipeline. Despite our effort to collect repositories with diverse code, our dataset for offline evaluation remains limited in terms of size and might not represent all the code well enough.
Online evaluation bias. Our study utilized anonymous data from Early Access Program (EAP) users to evaluate various code completion approaches via A/B testing. Although the results were statistically significant, they may not fully represent the broader user base of IntelliJ-based IDEs, as EAP participants typically have a deeper understanding and engagement with the IDE’s features, and their usage patterns may differ from average users.
@@
VII Conclusion
In this work, we shared our approach to implementing a multi-token code completion feature for IntelliJ-based IDEs, called Full Line Code Completion. The feature fills a niche of local-based multi-token code completion engines, being available for a wide range of users, including those with limited internet access and those with privacy concerns. It meets the speed and memory footprint restrictions, as well as fulfills several other important design principles. The resulting implementation has proven its efficiency during A/B testing on real users and was bundled into 2024.1 & 2024.2 releases of IntelliJ-based IDEs. Our ongoing and further work is focused on different languages and making sure that the feature performs well in all of them. We hope that our work can prove useful to both researchers and practitioners, and showcase the lifecycle of a complex research project in the real product.